OpenStudy (anonymous):

9^20 + 9^20 + 9^20? Please help me!

OpenStudy (anonymous):

so it should gives 3(9^20)= 3 x 9^20 = 3 x (3^2)^20 = 3 x 3^(40) = 3^41

OpenStudy (anonymous):

:D thank you!

OpenStudy (anonymous):

Khong co chi! :D

OpenStudy (anonymous):

hahaha

OpenStudy (anonymous):

lol haha

OpenStudy (anonymous):

are you viet?

OpenStudy (anonymous):

look like i can speak vietnamese, so im viet :P

OpenStudy (anonymous):

ahhh just checking

OpenStudy (anonymous):

don't worry, i got you anytime you need help with math ;)

OpenStudy (anonymous):

OH! Yay Thank-You ;D

OpenStudy (anonymous):

that's an incredibly larger number and the answer probably wouldn't fit the width of your computer screen

OpenStudy (anonymous):

LOL nice one :L